Grade 11Mechanics

Two particles of equal mass M,go round a circle of radius R under the action of mutual gravitational attraction.The speed of each particle is:

Spicer
8 Years agoGrade 11

1 Answer

Dhyaan
8 Years ago
As the two particles moving in circular motion the distance of separation is 2R. The gravitational force according to the Newton's law of gravitation is:
GM^2/(2R)^2=mv^2/R
Since mutual gravitational force is the only force providing centripetal acceleration
which implies  V is= ( GM / 4 R)^1/2